OpenStudy (anonymous):

Maya and Sally were the two finalists in a singing competition. The person with the most votes from the audience was chosen as the winner. Sally received 25% of the total votes and lost by 62,500 votes. No person in the audience was allowed to vote more than once. The equation below represents this situation, where y is the total number of votes cast and x is the number of votes Maya received. x - 62,500 = 0.25y If a total of 125,000 people voted, how many votes did Maya receive?

OpenStudy (anonymous):

i am confused is it like x = Maya, y = total, = 125,000 x - 62500 = 0.25(125,000) = 31,250 x - 62500 = 31,250 x = 62500 + 31250 = 93,750 votes for Maya i am not sure

OpenStudy (lgbasallote):

right again

OpenStudy (anonymous):

x = Maya, y = total, = 125,000 x - 62500 = 0.25(125,000) = 31,250 x - 62500 = 31,250 x = 62500 + 31250 = 93,750 votes for Maya

OpenStudy (anonymous):

Its right

OpenStudy (anonymous):

the problem that i am not always sure in my answer because i need to get full marks =)

OpenStudy (anonymous):

BE \[\huge\ CONFIDENT\]

OpenStudy (lgbasallote):

that's how i feel in quizzes to lol. they seem so easy that you feel nervous you're wrong, and it's a trick question

OpenStudy (anonymous):

thanks guys